If observed time = .1 h , rating = 110%, and shift allowances = 20%, what is the standard time.
Answer #1

Normal time = observed time * rating factor

Normal time = 0.1 * 110/100 = 0.11 hours

Standard time = Normal time / (1 – allowance factor)

Standard time = 0.11/ (1 – (20/100)) = 0.1375 hours

Standard time = 0.1375 hours or 8.25 minutes